PRESCRIBING SCALAR AND BOUNDARY MEAN CURVATURE ON THE THREE DIMENSIONAL HALF SPHERE

ZINDINE DJADLI ANDREA MALCHIODI MOHAMEDEN OULD AHMEDOU

Abstract. We consider the problem of prescribing the scalar curvature and the boundary mean curvature of the standard half three sphere, by deforming conformally its standard metric. Using blow up analysis techniques, we prove some existence and compactness results.